A satellite is in a circular orbit around the planet saturn. Derive the relationship between the period of the satellite, the mass of Saturn and the radius of orbit

1
Expert's answer
2021-11-29T11:40:28-0500

Explanations & Calculations


  • Using Newton's law of gravitation for the motion of the satellite around Saturn, this can be derived as follows (mass of Saturn = M, radius of the orbit = r, gravitational constant =G),

"\qquad\qquad
\begin{aligned}
\small G\frac{Mm}{r^2}&=\small mr\omega^2\\
\small G\frac{M}{r^2}&=\small r\Big(\frac{2\pi}{T}\Big)^2=r.\frac{4\pi^2}{T^2}\\
\small T^2&=\small 4\pi^2.\frac{r^3}{GM}\\
\small T&=\small 2\pi.\sqrt\frac{r^3}{GM}
\end{aligned}"